# Aditya Tanna, Pratinav Seth, Mohamed Bouadi Utsav Avaiya, Vinay Kumar Sankarapu Lexsi Labs, India & France

### v2

@@ 1  Introduction
Tabular data—structured collections of records (rows) and attributes (columns)—underpins a vast array of real-world applications, from healthcare and finance to industrial analytics and scientific research. These datasets form the backbone of modern decision-making systems: hospitals rely on patient records for diagnostic risk prediction, financial institutions assess creditworthiness and fraud using structured features, and scientific studies aggregate multi-sensor measurements to discover new insights. Yet despite this ubiquity, deep learning on tabular data has long lagged behind the breakthroughs achieved in domains such as natural language processing and computer vision.
@@
Recently, the advent of tabular foundation models (TFMs) has begun to reshape this landscape. These models extend the “foundation model” paradigm—pretraining on large, diverse corpora followed by downstream adaptation—to structured data domains. Models such as TabPFN [tabpfn, TabPFN2], TabICL [tabicl], OrionMSP [orionmsp2025], and OrionBiX [orionbix2025] leverage large-scale pre-training on synthetic or massive tabular datasets to provide generic, out-of-the-box learning capabilities.
For example, TabPFN demonstrates strong few-shot classification performance on small to medium datasets without additional gradient updates, approximating Bayesian posterior inference in a single forward pass. Meanwhile, TabICL extends this paradigm to larger datasets, scaling in-context learning to hundreds of thousands of samples, while the Orion family introduces efficient multi-scale and biaxial attention mechanisms that enhance scalability and accuracy. These developments mark a significant step toward “general-purpose” tabular learners that can adapt rapidly to new datasets with minimal supervision.
@@ 2.2  Tabular Foundation Models
TabPFN (Tabular Prior-data Fitted Network) represents one of the earliest and most influential TFMs. It employs a transformer architecture trained on vast collections of synthetic tabular datasets, effectively learning to approximate Bayesian inference across varying generative priors. This pre-training strategy enables TabPFN to achieve strong few-shot and zero-shot performance on small to medium datasets (typically up to $\sim$10,000 samples), offering competitive accuracy without explicit fine-tuning [tabpfn, TabPFN2].
Building on this foundation, TabICL extends the in-context learning paradigm to larger-scale tabular datasets by introducing a novel two-stage column-then-row attention mechanism. Through large-scale synthetic pre-training (up to $\sim$60K samples) and robust scaling to half a million samples at inference time, TabICL demonstrates that transformer-based ICL can challenge traditional tree ensembles even on large tabular domains [tabicl].
@@
Another direction, TabDPT, introduces denoising-based or self-supervised pre-training strategies that learn robust feature representations from both synthetic and real-world tables [tabdpt]. In parallel, other approaches explore the role of mixed synthetic priors, showing that a combination of causal, statistical, and randomized priors during pre-training can substantially improve downstream generalization and calibration performance across classification and regression tasks. Collectively, these developments illustrate that, with sufficient architectural bias and large-scale pre-training, transformer-based TFMs can achieve broad generalization on structured data—challenging long-held assumptions about the superiority of tree-based ensembles in tabular learning. Recent comprehensive surveys have catalogued the landscape of tabular foundation models [jiang2025representation, badaro2023transformers], highlighting their potential to transform tabular learning. Recent work has demonstrated that careful fine-tuning strategies can yield significant improvements for TFMs [rubachev2025finetuning]. Meanwhile, researchers have argued that tabular foundation models should be a research priority [breugel2024position], given their potential impact across domains.
@@
OrionMSP [orionmsp2025] introduces a transformer architecture for tabular in-context learning that combines multi-scale sparse attention—capturing both local and global feature dependencies efficiently—with a Perceiver-style latent memory that enables safe bidirectional communication between training and test samples. It uses column-wise embeddings via Set Transformers to model feature distributions and a split-masked causal attention mechanism for proper ICL reasoning. The design achieves near-linear attention scaling while maintaining high accuracy across diverse datasets, rivaling or outperforming TabICL and TabPFN—especially on heterogeneous, high-dimensional, and imbalanced datasets—thus demonstrating robust, efficient, and generalizable tabular learning without gradient updates.
Similarly, OrionBiX [orionbix2025] employs a biaxial attention mechanism for tabular in-context learning, enabling bidirectional context modeling across both rows and columns of tabular data. This dual-axis attention pattern allows the model to capture complex feature interactions and sample relationships simultaneously, providing a more comprehensive representation of tabular structure. The biaxial design facilitates effective information flow between training and test samples, enabling strong in-context learning performance across diverse tabular datasets while maintaining computational efficiency.
@@ 3.3  Supported Models
TabDPT Denoising Transformer Employs diffusion-based self-supervised pre-training
@@ 6  Experimental Results
We conduct a comprehensive evaluation of TabTune across multiple dimensions: predictive performance, probability calibration, fairness, and scalability. This section presents our experimental design and detailed results across diverse tabular learning scenarios.
@@ 6.1  Experimental Design
6.1.1  Benchmark Suites and Datasets.
Our experiments encompass three well-established benchmark suites—TALENT [talent] (181 automatically discovered datasets), OpenML-CC18 [openmlcc18] (72 curated datasets), and TabZilla [tabzilla] (36 heterogeneous tasks)—facilitating a systematic comparison across a broad spectrum of tabular learning settings. We further conduct domain-specific evaluations in high-impact areas such as medical and financial domains to assess the practical applicability of the methods. All experiments adhere to the standardized dataset splits defined by each benchmark, ensuring both reproducibility and fairness in comparison.
@@
To maintain consistency and fairness across model families, benchmark results are computed only on the common subset of datasets available for all evaluated models within each benchmark suite. This unified evaluation protocol ensures that performance rankings and metrics reflect genuine methodological differences rather than disparities in dataset coverage. In total, our evaluations span 155/181 datasets for TALENT, 27/36 datasets for TabZilla, and 63/72 datasets for OpenML-CC18. Some datasets were excluded due to out-of-memory (OOM) errors even on H200 GPUs or CUDA-related issues, particularly affecting TabPFN-based models.
It is important to clarify how mean rank values are derived. Within each benchmark suite, models are ranked by accuracy on every dataset (lower rank = better performance), and these per-dataset ranks are averaged to obtain the overall mean rank. Thus, a lower mean rank indicates stronger and more consistent performance across datasets, rather than the highest score on any single task. While absolute metrics (accuracy, F1) reflect peak task-level performance, mean rank provides a normalized measure of cross-dataset generalization consistency.
Detailed dataset statistics are provided in Appendix 11.
6.1.2  Models and Adaptation Strategies.
@@
We evaluate seven recent tabular foundation models—TabPFN, TabICL, OrionMSP, OrionBiX, Mitra, ContextTab, and TabDPT—under multiple adaptation paradigms, including zero-shot inference, meta-learning fine-tuning, supervised fine-tuning, parameter-efficient supervised fine-tuning, and parameter-efficient meta-learning. In addition, we include established traditional baselines using autogloun [erickson2020autogluon] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.
6.1.3  Hardware Configuration.
Experiments are executed on NVIDIA L40S GPUs, with H200 GPUs used for memory-intensive cases. This infrastructure ensures consistent execution across all experiments while handling the computational demands of large transformer-based models.
6.1.4  Evaluation Metrics.
Our assessment covers four complementary dimensions:
Performance:
@@
Classification Accuracy (ACC), AUC-ROC, and weighted F1-score (F1) quantify predictive capability across standard benchmark datasets (TALENT, OpenML-CC18, TabZilla) covering diverse dataset characteristics including small and large sample sizes, narrow and wide feature spaces, and balanced and imbalanced class distributions.
Scalability:
@@
Analysis of performance variation across dataset sizes, feature dimensionality, and class imbalance provides practical guidance for model selection. This analysis uses the same benchmark datasets as performance evaluation, aggregated and summarized across these dimensions to reveal scalability patterns.
Calibration:
@@
We evaluate how fine-tuning impacts calibration metrics across different models and adaptation strategies, measuring the alignment between predicted probabilities and actual outcomes, which is critical for decision-making systems. Our evaluation uses three standard metrics:
  • Expected Calibration Error (ECE), which measures the average deviation between predicted confidence and observed accuracy across binned predictions;
  • Maximum Calibration Error (MCE), which captures the worst-case calibration gap, identifying regions of systematic miscalibration; and
  • Brier score, which quantifies both calibration and refinement by penalizing squared deviations between predicted probabilities and binary outcomes.
Fairness:
We assess equitable treatment across demographic subgroups using established fairness-oriented datasets with sensitive attributes (gender, age, race, income group). We measure three complementary fairness criteria that capture different notions of equitable treatment:
  • Statistical Parity Difference (SPD) quantifies the absolute difference in positive prediction rates across groups, testing whether the model provides equal opportunity independent of sensitive attributes. A model satisfies demographic parity when SPD approaches zero, indicating that positive predictions are distributed proportionally across demographic groups.
  • Equalized Odds Difference (EOD) measures disparities in both true positive and false positive rates across groups, ensuring consistent error profiles across subgroups. This criterion is more stringent than statistical parity, requiring not only proportional positive predictions but also equal accuracy within each group.
  • Equalized Opportunity Difference (EOpD) focuses specifically on true positive rate parity, ensuring that qualified individuals receive equal treatment regardless of group membership. This metric is particularly relevant in settings where false negatives have severe consequences, such as loan approvals or hiring decisions.
6.1.5  Analysis Focus :
@@
A central goal of our evaluation is to analyze how different adaptation strategies—zero-shot inference, meta-learning fine-tuning, and supervised fine-tuning—affect model behavior across multiple dimensions. We examine whether fine-tuning improves performance, influences calibration, and impacts fairness across domains. This comparative analysis offers practical guidance on selecting appropriate adaptation strategies for tabular foundation models in different application contexts.
@@ 6.2  Performance Evaluation
6.2.1  Overall Leaderboards
Table 3 presents comprehensive results across the TALENT, OpenML-CC18, and TabZilla benchmark suites, reporting mean rank, classification accuracy (ACC), and weighted F1-score (F1) for each model and adaptation strategy. The aggregated findings reveal clear and consistent trends in how adaptation mechanisms influence predictive performance and cross-dataset generalization.
Baselines and Zero-Shot Inference.
Classical Machine Learning methods remain strong reference baselines, attaining mean accuracies between 0.833 and 0.861 with aggregated ranks around 6.0. In contrast, pretrained tabular foundation models (TFMs) exhibit markedly stronger generalization even without task-specific training. Among these, OrionMSP achieves the best overall zero-shot rank of 3.58, reaching (ACC = 0.8461/F1 = 0.8360) on TALENT, (0.8722/0.8676) on OpenML-CC18, and (0.8821/0.8786) on TabZilla. TabPFN follows closely (overall rank - 4.61) with (ACC = 0.8514/F1 = 0.8412) on TALENT and up to (0.8752/0.8716) on TabZilla. TabDPT (Rank - 5.42) attains (ACC = 0.8408 / F1 = 0.8318) on TALENT and (0.8814/0.8775) on TabZilla, while Mitra (Rank - 11.77, ACC < 0.40) and ContextTab (Rank - 9.70) perform substantially worse. Overall, zero-shot TFMs exceed traditional machine learning models by roughly 2–4 percentage points in accuracy and three to four rank positions, demonstrating the strength of pretrained priors.
@@ 6.2.1  Overall Leaderboards
XGBoost 6.70 6.02 0.8403 0.8360 5.89 0.8558 0.8537 6.07 0.8612 0.8326
CatBoost 6.43 5.57 0.8336 0.8259 6.25 0.8588 0.8520 7.13 0.8579 0.8384
Random Forest 7.38 6.15 0.8285 0.8209 6.36 0.8547 0.8497 8.42 0.8358 0.8399
LightGBM 6.78 6.11 0.8331 0.8245 6.18 0.8581 0.8493 5.25 0.8618 0.8211
TabICL 4.96 4.09 0.8471 0.8379 4.69 0.8667 0.8623 5.89 0.8734 0.8698
OrionBiX 5.37 4.59 0.8346 0.8260 4.98 0.8653 0.8596 4.89 0.8728 0.8628
OrionMSP 3.58 3.26 0.8461 0.8360 4.12 0.8722 0.8676 3.84 0.8821 0.8786
TabPFN 4.61 3.72 0.8514 0.8412 4.76 0.8714 0.8663 4.86 0.8752 0.8716
Mitra 11.77 10.38 0.3921 0.2868 10.52 0.3614 0.2522 11.21 0.3152 0.1830
ContextTab 9.70 9.84 0.5474 0.4596 6.28 0.8639 0.8581 7.13 0.8389 0.8334
TabDPT 5.42 5.19 0.8408 0.8318 4.64 0.8672 0.8625 3.94 0.8814 0.8775
@@
Meta-Learning Fine-Tuning.

### v3

@@ 1  Introduction
Tabular data i.e. structured collections of records (rows) and attributes (columns)—underpins a vast array of real-world applications, from healthcare and finance to industrial analytics and scientific research. These datasets form the backbone of modern decision-making systems: hospitals rely on patient records for diagnostic risk prediction, financial institutions assess creditworthiness and fraud using structured features, and scientific studies aggregate multi-sensor measurements to discover new insights. Yet despite this ubiquity, deep learning on tabular data has long lagged behind the breakthroughs achieved in domains such as natural language processing and computer vision.
@@
Recently, the advent of tabular foundation models (TFMs) has begun to reshape this landscape. These models extend the “foundation model” paradigm—pretraining on large, diverse corpora followed by downstream adaptation—to structured data domains. Models such as TabPFN [tabpfn, TabPFN2, tabpfn25], TabICL [tabicl], ContextTAB [contexttab], OrionMSP [orionmsp2025], and OrionBiX [orionbix2025] leverage large-scale pre-training on synthetic or massive tabular datasets to provide generic, out-of-the-box learning capabilities. These developments mark a significant step toward “general-purpose” tabular learners that can adapt rapidly to new datasets with minimal supervision.
@@ 2.2  Tabular Foundation Models
TabPFN (Tabular Prior-data Fitted Network) represents one of the earliest and most influential TFMs. It employs a transformer architecture trained on vast collections of synthetic tabular datasets, effectively learning to approximate Bayesian inference across varying generative priors. This pre-training strategy enables TabPFN to achieve strong few-shot and zero-shot performance on small to medium datasets (typically up to $\sim$10,000 samples), offering competitive accuracy without explicit fine-tuning [tabpfn, TabPFN2, tabpfn25]. Building on this foundation, TabICL extends the in-context learning paradigm to larger-scale tabular datasets by introducing a novel two-stage column-then-row attention mechanism. Through large-scale synthetic pre-training (up to $\sim$60K samples) and robust scaling to half a million samples at inference time, TabICL demonstrates that transformer-based ICL can challenge traditional tree ensembles even on large tabular domains [tabicl]. Another line of work, TabDPT, introduces an in-context learning (ICL)-based tabular foundation model trained with denoising-style, self-supervised objectives on real-world tables [tabdpt]. In parallel, other approaches explore the role of mixed synthetic priors, showing that combining causal, statistical, and randomized priors during pre-training can substantially improve downstream generalization and calibration performance across classification and regression tasks.
@@
Collectively, these developments illustrate that, with sufficient architectural bias and large-scale pre-training, transformer-based TFMs can achieve broad generalization on structured data—challenging long-held assumptions about the superiority of tree-based ensembles in tabular learning. Recent comprehensive surveys have catalogued the landscape of tabular foundation models [jiang2025representation, badaro2023transformers], highlighting their potential to transform tabular learning. Recent work has demonstrated that careful fine-tuning strategies can yield significant improvements for TFMs [rubachev2025finetuning]. Meanwhile, researchers have argued that tabular foundation models should be a research priority [breugel2024position], given their potential impact across domains.
@@
OrionMSP [orionmsp2025] introduces a transformer architecture for tabular in-context learning that combines multi-scale sparse attention—capturing both local and global feature dependencies efficiently—with a Perceiver-style latent memory that enables safe bidirectional communication between training and test samples. It uses column-wise embeddings via Set Transformers to model feature distributions and a split-masked causal attention mechanism for proper ICL reasoning. The design achieves near-linear attention scaling while maintaining high accuracy across diverse datasets, rivaling or outperforming TabICL and TabPFN—especially on heterogeneous, high-dimensional, and imbalanced datasets—thus demonstrating robust, efficient, and generalizable tabular learning without gradient updates. Similarly, OrionBiX [orionbix2025] employs a biaxial attention mechanism for tabular in-context learning, enabling bidirectional context modeling across both rows and columns of tabular data. This dual-axis attention pattern allows the model to capture complex feature interactions and sample relationships simultaneously, providing a more comprehensive representation of tabular structure. The biaxial design facilitates effective information flow between training and test samples, enabling strong in-context learning performance across diverse tabular datasets while maintaining computational efficiency.
@@ 3.3  Supported Models
TabDPT In-Context Learning Employs diffusion-based self-supervised pre-training
@@ 4.3  Evaluation Utilities
4.4  Installation
TabTune is designed to be easily integrated into existing Python environments. The library is open-source and available on the Python Package Index (PyPI), ensuring straightforward installation for practitioners.
To install the latest stable release, users can execute the following command:
1pip install tabtune
Alternatively, for the latest development version, the package can be installed directly from the source repository:
1git clone https://github.com/Lexsi-Labs/TabTune.git
2cd TabTune
3pip install -r requirements.txt
4pip install -e .
@@ 6  Experimental Results
We conduct evaluation of TabTune across multiple dimensions: predictive performance, probability calibration, fairness, and scalability. This section presents our experimental design and detailed results across diverse scenarios.
@@ 6.1  Experimental Design
Benchmark Suites and Datasets : Our experiments encompass three well-established benchmark suites i.e. TALENT [talent] (181 datasets), OpenML-CC18 [openmlcc18] (72 datasets), and TabZilla [tabzilla] (36 datasets). Facilitating a systematic comparison across a broad spectrum of tabular learning settings. We further conduct domain-specific evaluations in high-impact areas such as medical and financial domains to assess the practical applicability of the methods. All experiments adhere to the standardized dataset splits defined by each benchmark, ensuring both reproducibility and fairness in comparison. To maintain consistency and fairness across model families, benchmark results are computed only on the common subset of datasets available for all evaluated models within each benchmark suite. This unified evaluation protocol ensures that performance rankings and metrics reflect genuine methodological differences rather than disparities in dataset coverage. In total, our evaluations span 155/181 datasets for TALENT, 27/36 datasets for TabZilla, and 63/72 datasets for OpenML-CC18. Some datasets were excluded due to out-of-memory (OOM) errors even on H200 GPUs or CUDA-related issues, particularly affecting TabPFN-based models.
@@
Models and Adaptation Strategies: We evaluate recent tabular foundation models—TabPFN, TabICL, OrionMSP, OrionBiX, Mitra, ContextTab, and TabDPT—under multiple adaptation paradigms, including zero-shot inference, meta-learning fine-tuning, supervised fine-tuning, parameter-efficient supervised fine-tuning, and parameter-efficient meta-learning. In addition, we include established traditional baselines using AutoGluon [erickson2020autogluon] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.
Evaluation Metrics : Our assessment covers four complementary dimensions:
@@
Performance: Classification Accuracy (ACC), AUC-ROC, and weighted F1-score (F1) quantify predictive capability across standard benchmark datasets (TALENT, OpenML-CC18, TabZilla) covering diverse dataset characteristics including small and large sample sizes, narrow and wide feature spaces, and balanced and imbalanced class distributions. It is important to clarify how mean rank values are derived. Within each benchmark suite, models are ranked by accuracy on every dataset (lower rank = better performance), and these per-dataset ranks are averaged to obtain the overall mean rank. Thus, a lower mean rank indicates stronger and more consistent performance across datasets, rather than the highest score on any single task. While absolute metrics (accuracy, F1) reflect peak task-level performance, mean rank provides a normalized measure of cross-dataset generalization consistency.
@@
Scalability: Analysis of performance variation across dataset sizes, feature dimensionality, and class imbalance provides practical guidance for model selection. This analysis uses the same benchmark datasets as performance evaluation, aggregated and summarized across these dimensions to reveal scalability patterns.
@@
Calibration: We evaluate how fine-tuning impacts calibration metrics across different models and adaptation strategies, measuring the alignment between predicted probabilities and actual outcomes, which is critical for decision-making systems. Our evaluation uses three metrics: (I) Expected Calibration Error (ECE), which measures the average deviation between predicted confidence and observed accuracy across binned predictions; (II) Maximum Calibration Error (MCE), which captures the worst-case calibration gap, identifying regions of systematic miscalibration; and (III) Brier score, which quantifies both calibration and refinement by penalizing squared deviations between predicted probabilities and binary outcomes.
Fairness: We assess equitable treatment across demographic subgroups using established fairness-oriented datasets with sensitive attributes (gender, race). We measure three fairness criteria that capture different notions of equitable treatment: (I) Statistical Parity Difference (SPD) quantifies the absolute difference in positive prediction rates across groups, testing whether the model provides equal opportunity independent of sensitive attributes. A model satisfies demographic parity when SPD approaches zero, indicating that positive predictions are distributed proportionally across demographic groups. (II) Equalized Odds Difference (EOD) measures disparities in both true positive and false positive rates across groups, ensuring consistent error profiles across subgroups. This criterion is more stringent than statistical parity, requiring not only proportional positive predictions but also equal accuracy within each group. (III) Equalized Opportunity Difference (EOpD) focuses specifically on true positive rate parity, ensuring that qualified individuals receive equal treatment regardless of group membership. This metric is particularly relevant in settings where false negatives have severe consequences, such as loan approvals or hiring decisions.
@@
Analysis Focus : A central goal of our evaluation is to analyze how different adaptation strategies—zero-shot inference, meta-learning fine-tuning, and supervised fine-tuning—affect model behavior across multiple dimensions. We examine whether fine-tuning improves performance, influences calibration, and impacts fairness across domains. This comparative analysis offers practical guidance on selecting appropriate adaptation strategies for tabular foundation models in different application contexts. Detailed dataset statistics and information refer to Appendix 10.
All Experiments are executed on NVIDIA L40S (48Gb VRAM), with H200 (141Gb VRAM) used for memory-intensive cases. This ensures consistent execution across all experiments while handling the computational demands of models.
@@ 6.2  Performance Evaluation
XGBoost 7.94 7.43 0.8395 0.8351 6.57 0.8558 0.8537 6.82 0.8612 0.8326
CatBoost 7.69 6.90 0.8329 0.8252 6.98 0.8588 0.8520 6.70 0.8579 0.8384
Random Forest 8.71 7.56 0.8278 0.8201 7.10 0.8547 0.8497 9.27 0.8358 0.8399
LightGBM 8.03 7.51 0.8324 0.8238 6.89 0.8581 0.8493 5.80 0.8618 0.8211
TabICL 5.92 5.09 0.8465 0.8372 5.30 0.8667 0.8623 6.65 0.8734 0.8698
OrionBix 6.52 5.87 0.8339 0.8253 5.63 0.8653 0.8596 5.47 0.8728 0.8628
OrionMSP 4.61 4.06 0.8455 0.8353 4.58 0.8722 0.8676 4.36 0.8821 0.8786
TabPFN-2.0 5.45 4.48 0.8508 0.8406 5.48 0.8714 0.8663 5.65 0.8752 0.8716
TabPFN-2.5 4.13 3.62 0.8540 0.8440 4.36 0.8726 0.8672 3.97 0.8824 0.8791
Mitra 13.07 11.63 0.3935 0.2884 11.50 0.3614 0.2522 12.21 0.3152 0.1830
ContextTab 7.75 6.52 0.8373 0.8269 7.01 0.8639 0.8581 8.02 0.8389 0.8334
TabDPT 6.45 6.33 0.8402 0.8311 5.18 0.8672 0.8625 4.57 0.8814 0.8775
@@
6.2.1  Overall Leaderboards
Table 3 presents comprehensive results across the TALENT, OpenML-CC18, and TabZilla benchmark suites, reporting mean rank, accuracy (ACC), and weighted F1-score (F1) for each model and adaptation strategy. The aggregated findings reveal clear and consistent trends in how adaptation mechanisms influence predictive performance and cross-dataset generalization.
Baselines and Zero-Shot Inference : Classical Machine Learning methods remain strong reference baselines, attaining mean accuracies between 0.833 and 0.861 with aggregated ranks around 6.0. In contrast, pretrained tabular foundation models (TFMs) exhibit markedly stronger generalization even without task-specific training.
Among pretrained TFMs, TabPFN-2.5 now achieves the best overall zero-shot rank (4.13), with accuracies of (ACC=0.8540/F1=0.8440) on TALENT, (ACC=0.8726/F1=0.8672) on OpenML-CC18, and (ACC=0.8824/F1=0.8791) on TabZilla. OrionMSP follows closely (overall rank 4.61), reaching (ACC=0.8455 / 0.8353) on TALENT, (ACC=0.8722/F1=0.8676) on OpenML-CC18, and (ACC=0.8821/F1=0.8786) on TabZilla. TabICL and TabDPT remain competitive (ranks 5.92 and 6.45, ACC $\approx$ 0.84–0.88), while Mitra and ContextTab lag behind. Overall, zero-shot TFMs exceed traditional machine-learning baselines by roughly 2–4 percentage points in accuracy and three to four rank positions, demonstrating the strength of pretrained priors.